# Forecasting Fold Bifurcations through Physics-Informed Convolutional Neural Networks

### v1

@@ Abstract
This study proposes a physics-informed convolutional neural network (CNN) for identifying dynamical systems’ time series near a fold bifurcation. The peculiarity of this work is that the CNN is trained with a relatively small amount of data and on a single, very simple system. In contrast, the CNN is validated on much more complicated systems. A similar task requires significant extrapolation capabilities, which are obtained by exploiting physics-based information. Physics-based information is provided through a specific pre-processing of the input data, consisting mostly of a transformation into polar coordinates, normalization, transformation into the logarithmic scale, and filtering through a moving mean. The results illustrate that such data pre-processing enables the CNN to grasp the important features related to approaching a fold bifurcation, namely, the trend of the oscillation amplitude, and neglect other characteristics that are not particularly relevant, such as the vibration frequency. The developed CNN was able to correctly classify trajectories near a fold for a mass-on-moving-belt system, a van der Pol-Duffing oscillator with an attached tuned mass damper, and a pitch-and-plunge wing profile. The results obtained pave the way for the development of similar CNNs effective in real-life applications.
@@
Physics-informed neural network, convolutional neural network, fold bifurcation, saddle-node bifurcation, bifurcation prediction
@@ 1 Introduction
Fold, or saddle-node, bifurcations are a type of bifurcation related to the merging of a saddle and a node steady-state solutions; hence the name saddle-node bifurcation kuznetsov1998elements . The saddle and node solutions can have different natures, typically equilibrium points or periodic solutions. In both cases, if a point of each solution is represented in a plane for variations of a parameter, the branch of the two solutions generally assumes a folded shape, as shown in Fig. 1, from which the other name fold bifurcation is derived. Accordingly, fold bifurcations mark the appearance of two steady-state solutions, one stable and one unstable. In this study, we refer uniquely to fold bifurcations of periodic solutions.
@@
Fold bifurcations have an important role in engineering practice. Let us consider the bifurcation scenario in Fig. 1 and assume that the upper branch of stable steady-state solutions represents the working condition of a system. If the bifurcation parameter decreases below the value of the fold bifurcation, the desired steady-state ceases to exist, leading the system to suddenly diverge from its state, usually catastrophically. Accordingly, this scenario is a mathematical catastrophe arnold1981singularity .
@@
On the other hand, if the desired solution is not directly related to the fold (for instance, the trivial equilibrium in Fig. 1), then the fold does not affect the system’s local dynamics around that equilibrium. In fact, this solution does not change its stability properties when the bifurcation parameter crosses the value of the fold bifurcation. This scenario suggests that the fold bifurcation can be neglected in this case, which is often done in practice. However, the fold can have an important effect on the system’s global dynamics. Let us consider the scenario in Fig. 1, assuming that no other solution exists; in regions A and B, before the fold, the trivial solution is globally stable, while after the fold, it is only locally stable. This means that, in the locally stable region, while a small perturbation will only generate a transient effect, a large enough perturbation might lead the system to converge towards the large amplitude solutions. In practice, this scenario is particularly elusive; in fact, even after extensively testing the system, initial conditions leading to the upper branch of solutions might never be encountered, leaving the fold bifurcation undetected. However, during regular operations, it might be enough to have such perturbation only once to cause an accident. This phenomenon is usually referred to as the dynamical integrity of a stable solution, which is unbounded only for globally stable solutions lenci2019global .
@@
Several systems of engineering relevance suffer from limited dynamical integrity and local stability. Examples include flutter instability ghadami2016bifurcation ; nitti2021spatially ; habib2021dynamical , machining processes dombovari2019experimental ; szaksz2024dynamical , brake squeal papangelo2017subcritical ; hu2020friction , robot control habib2013nonlinear ; habib2022bistability , wheel shimmy horvath2022stability ; habib2023towed , pressure relief valves kadar2023nonlinear , turbulent flows cherubini2015nonlinear ; kerswell2018nonlinear , electric blackouts gajduk2014stability ; ren2015early , human brain epilepsy suffczynski2004dynamics ; lytton2008computer , human balance zakynthinaki2010modeling ; smith2017basins and prey-predator ecosystems saleh2022basins ; garai2023coexistence , to name a few.
@@
Quite often, regions of local stability are related to subcritical bifurcations. For example, subcritical Andronov-Hopf bifurcations generate a branch of unstable periodic solutions that coexists with the stable one; subcritical Neimark-Sacker bifurcations generate a similar scenario kuznetsov1998elements , but with quasiperiodic solutions. However, in some cases, limited dynamical integrity is due to isolated branches of periodic solutions, which makes their detection even harder habib2018isolated . These are encountered in several systems, such as chemical reactors uppal1976classification , traffic models martinovich2023nonlinear , bridges under wind excitation zulli2012bifurcation , and shimmy of rigid wheels takacs2008isolated , for example.
@@
Several methods exist to identify and predict fold bifurcations. Analytical methods typically provide an approximate solution to the system dynamics, usually in the form of a system of algebraic equations. This applies to the multiple-scale nayfeh2008nonlinear , averaging verhulst2023toolbox , or harmonic balance methods, to mention a few. Once a system of algebraic equations for the system’s solution is available, a fold can be found either by plotting the whole solutions in a given parameter range or by studying the system singularities through implicit derivatives of the system of algebraic equations golubitsky1985singularities . Clearly, implementing these methods requires knowledge of the system’s equations of motion.
@@
Numerical methods to identify fold bifurcations mostly consist of pseudo-arclength algorithms coupled with shooting techniques. By varying the bifurcation parameter, one of the two branches of periodic solutions is continued until the fold is reached nayfeh2008applied . Alternatively, through direct numerical integration of the equations of motion and by performing a sweep-up and a sweep-down of the bifurcation parameter, the fold is defined by a sudden jump-down and settling of the periodic solution onto an equilibrium (for a bifurcation scenario similar to the one depicted in Fig. 1). Although these methods typically exploit the system’s equations of motion, they can also be implemented through a black-box model of the system.
@@
Similarly, experimental methods for fold bifurcation identification usually implement a sweep-up and a sweep-down of the bifurcation parameter while the system operates. The fold is identified by a sudden jump-down of a periodic solution. Additionally, control-based continuation techniques enable tracking unstable solutions until a fold is reached or even track folds in the parameter space renson2019numerical .
@@
Several methods to predict the presence of fold bifurcations directly from time series, without actually reaching it, were proposed in the literature. Suppose the system is assumed on the branch of periodic solutions leading to the fold. In that case, these methods usually exploit early-warning signals related to statistical properties of the system’s response to small perturbations liu2015identifying ; ren2015early . These methods can recognize if a system is close to a fold (in this context, usually referred to as tipping point lenton2020tipping or critical transition scheffer2009early ; lade2012early ), possibly enabling operators to control the system and avoid reaching the fold, which would be a non-reversible transition.
@@
If the system in operation is on the branch of solutions not directly related to the fold, small perturbations can hardly be used to predict the fold; in fact, the fold has no local effect on the solution. However, large perturbations can still be used to forecast the fold without actually reaching it. Epureanu and coworkers developed similar methods in a series of papers lim2011forecasting ; ghadami2016bifurcation ; chen2018forecasting ; the methods they developed allow one to predict a full bifurcation diagram from the analysis of the displacement decay of a few trajectories obtained in the pre-bifurcation scenario. A similar technique, specifically targeting fold bifurcations, was recently developed in habib2023predicting and tested experimentally in kadar2024model .
@@
In this study, we try to perform a similar task utilizing machine learning. In particular, we utilize physics-informed convolutional neural networks (NNs). NNs have been long used to analyze time series and identify anomalies. Examples include but are not limited to, human activity recognition ordonez2016deep ; xu2019innohar , electrocardiogram classification zihlmann2017convolutional ; hannun2019cardiologist , speech recognition hinton2012deep ; nassif2019speech , natural language processing goldberg2016primer ; goldberg2022neural . NNs have the ability to handle long-term dependencies and temporal context yu2019review . They can recognize patterns and are good at performing classifications li2021survey . All these properties make them a potentially useful tool for identifying if a time series is near a fold bifurcation in the parameter space. Many studies demonstrated the ability of NNs to classify and process time series related to dynamical systems. For example, they are used for chatter detection in machining lamraoui2015chatter ; rahimi2021line , chaos identification boulle2020classification , fault detection and diagnosis in mechanical systems chen2021data ; junior2022fault , anomaly detection li2019deep , and many other similar tasks.
@@
Although NNs are excellent interpolators, being able to deal with large inputs, they are generally unable to extrapolate results, meaning that if an input is outside of the range of the input used for training, a NN tends to provide a wrong output zhu2023reliable . One method to overcome, or at least limit, this problem is to design and train the network, taking into account the physics of the problem as well, and not only data; in other words, by providing the network with some physics-based information. Such NNs are usually referred to as physics-informed NNs. The main idea behind them is to introduce an appropriate inductive, learning, or observational bias that can steer the learning process towards identifying physically consistent solutions karniadakis2021physics .
@@
Inductive bias consists of tailoring the NN architecture in such a way that it implicitly satisfies some physical laws; this approach is usually implemented through some sort of invariant symmetries and rotations; this is done, for instance, by convolutional NNs lecun1995convolutional , providing excellent results in computer vision and many other fields.
@@
Learning bias is usually provided through the loss function, which also includes a penalty if a given mathematical function is not satisfied. This mathematical function can have any desired form, and it should represent a physical principle that we want the NN to approximately satisfy, such as, for example, conservation of mass or energy in a mechanical system raissi2019physics .
@@
Observational bias consists of providing the network with data embodying the underlying physics of the problem or using specifically designed data augmentation procedures. For instance, an observational bias can be obtained by conveniently balancing experimental and synthetic data. This method was successfully applied in a variety of fields, such as, for example, fracture propagation chen2023physics .
@@
In this work, we investigate the ability of a physics-informed convolutional NN to identify trajectories that are close to a fold bifurcation. The NN is physics-informed through an observational bias obtained through an appropriate input normalization, which hides some signal properties (the frequency) and highlights other properties (the oscillation amplitude trend). Several studies in the literature have already approached the same problem through machine learning bury2021deep ; patel2023using ; however, up to the authors’ knowledge, it is the first time that the prediction is performed from trajectories not converging to the branch related to the fold itself. Our results show that such an NN can classify such trajectory with excellent accuracy and possesses remarkable extrapolation properties.
@@ 2 Objective
This study aims to train a convolutional NN (CNN) to identify if a time series belongs to a parameter set near a fold bifurcation. This objective is formalized in a classification task, as detailed below. The developed CNNs will be trained to distinguish between trajectories referring to a point in the parameter space that is far from a fold (region A in Fig. 1), close to a fold (region B), or after a fold (region C). Considering that the fold bifurcation generates two branches of periodic solutions, regions of the parameter space where the branches of periodic solutions do not exist are interpreted as before the fold (either far from it or close to it). In contrast, regions where they exist are considered as after.
We consider trajectories that are always initiated from relatively large initial conditions, such that they capture the transient dynamics related to the fold (although it might not be the case, as illustrated in habib2023predicting ). Other methods for forecasting fold bifurcations have, in fact, the same requirement lim2011forecasting ; habib2023predicting .
@@
We notice that, while the distinction between regions B and C (close and after the fold) is precisely defined by the bifurcation, the boundary between trajectories far and close to the fold is completely arbitrary. To overcome this problem, trajectories used to train the CNNs of the two groups are very well separated, as illustrated below. Then, the trained CNNs are tested on a large range of values of the bifurcation parameter in order to evaluate their capabilities.
@@ 3.1 Oscillator with nonlinear damping
For $c_{1}>0$, the trivial solution is always stable. For $c_{3}=0$, the system is linear, while for $c_{3}\neq 0$ the damping has a nonlinear characteristic. If $c_{3}>0$, for increasing oscillation amplitudes, first damping decreases and then increases because of the cubic and fifth order terms. As thoroughly investigated in habib2018isolated , for $c_{3}=c_{3}^{*}:=40c_{1}/9$ the system presents a fold bifurcation, which generates two branches of periodic solutions existing for $c_{3}>c_{3}^{*}$. Increasing $c_{3}$, the branch of unstable periodic solutions approaches the trivial solution, reducing its dynamical integrity. However, as far as $c_{1}>0$, they never merge. The corresponding bifurcation diagram is illustrated in Fig. 1(a).
@@
Figure 2: Bifurcation diagrams of the investigated systems; solid lines: stable solutions, dashed lines: unstable solutions; red dots: fold bifurcations. (a) System in Eq. (1) for $c_{1}=0.5$; (b) system in Eq. (2); (c) system in Eq. (5) for the parameter values $r=0.05$, $\gamma=0.97$, $\mu_{2}=0.12$ and $\alpha=0.3$; (d) system in Eq. (7).

### v2

# Fold Bifurcation Identification through Scientific Machine Learning

@@ Abstract
This study employs scientific machine learning to identify transient time series of dynamical systems near a fold bifurcation of periodic solutions. The unique aspect of this work is that a convolutional neural network (CNN) is trained with a relatively small amount of data and on a single, very simple system, yet it is tested on much more complicated systems. This task requires strong generalization capabilities, which are achieved by incorporating physics-based information. This information is provided through a specific pre-processing of the input data, which includes transformation into polar coordinates, normalization, transformation into the logarithmic scale, and filtering through a moving mean. The results demonstrate that such data pre-processing enables the CNN to grasp the important features related to transient time-series near a fold bifurcation, namely, the trend of the oscillation amplitude, and disregard other characteristics that are not particularly relevant, such as the vibration frequency. The developed CNN was able to correctly classify transient trajectories near a fold for a mass-on-moving-belt system, a van der Pol-Duffing oscillator with an attached tuned mass damper, and a pitch-and-plunge wing profile. The results contribute to the progress towards the development of similar CNNs effective in real-life applications such as safety monitoring of dynamical systems.
@@
Scientific machine learning , physics-informed neural network , convolutional neural network , fold bifurcation , saddle-node bifurcation , bifurcation prediction
journal: Physica D: Nonlinear Phenomena
@@ 1 Introduction
Fold, or saddle-node, bifurcations are a type of bifurcation related to the merging of a saddle and a node steady-state solutions; hence the name saddle-node bifurcation [1]. The saddle and node solutions can have different natures, typically equilibrium points or periodic solutions. In both cases, if a point of each solution is represented in a plane for variations of a parameter, the branch of the two solutions generally assumes a folded shape, as shown in Fig. 1, from which the other name fold bifurcation is derived. Accordingly, fold bifurcations mark the appearance of two steady-state solutions, one stable and one unstable. In this study, we refer uniquely to fold bifurcations of periodic solutions.
@@
Fold bifurcations have an important role in engineering practice. Let us consider the bifurcation scenario in Fig. 1 and assume that the upper branch of stable steady-state solutions represents the working condition of a system [2, 3]. If the bifurcation parameter decreases below the value of the fold bifurcation, the desired steady-state ceases to exist, leading the system to suddenly diverge from its state, usually catastrophically. Accordingly, this scenario is a mathematical catastrophe [4].
@@
On the other hand, if the desired solution is not directly related to the fold (for instance, the trivial equilibrium in Fig. 1), then the fold does not affect the system’s local dynamics around that equilibrium. In fact, this solution does not change its stability properties when the bifurcation parameter crosses the value of the fold bifurcation. This scenario suggests that the fold bifurcation can be neglected in this case, which is often done in practice. However, the fold can have an important effect on the system’s global dynamics. Let us consider the scenario in Fig. 1, assuming that no other solution exists; in regions A and B, before the fold, the trivial solution is globally stable, while after the fold, it is only locally stable. This means that, in the locally stable region, while a small perturbation will only generate a transient effect, a large enough perturbation might lead the system to converge towards the large amplitude solutions. In practice, this scenario is particularly elusive; in fact, even after extensively testing the system, initial conditions leading to the upper branch of solutions might never be encountered, leaving the fold bifurcation undetected. However, during regular operations, it might be enough to have such perturbation only once to cause an accident. This phenomenon is usually referred to as the dynamical integrity of a stable solution, which is unbounded only for globally stable solutions [5].
@@
Several systems of engineering relevance suffer from limited dynamical integrity and local stability. Examples include flutter instability [6, 7, 8], machining processes [9, 10], brake squeal [11, 12], robot control [13, 14], wheel shimmy [15, 16], pressure relief valves [17], turbulent flows [18, 19], electric blackouts [20, 21], human brain epilepsy [22, 23], human balance [24, 25] and prey-predator ecosystems [26, 27], to name a few.
@@
Quite often, regions of local stability are related to subcritical bifurcations. For example, subcritical Andronov-Hopf bifurcations generate a branch of unstable periodic solutions that coexists with the stable one; subcritical Neimark-Sacker bifurcations generate a similar scenario [1], but with quasiperiodic solutions. However, in some cases, limited dynamical integrity is due to isolated branches of periodic limit cycles, which makes their detection even harder [28]. These are encountered in several systems, such as chemical reactors [29], traffic models [30], bridges under wind excitation [31], and shimmy of rigid wheels [32], for example.
@@
Several methods exist to identify and predict fold bifurcations. Analytical methods typically provide an approximate solution to the system dynamics, usually in the form of a system of algebraic equations. This applies to the multiple-scale [33], averaging [34], or harmonic balance methods, to mention a few. Once a system of algebraic equations for the system’s solution is available, a fold can be found either by plotting the whole solutions in a given parameter range or by studying the system singularities through implicit derivatives of the system of algebraic equations [35]. Clearly, implementing these methods requires knowledge of the system’s equations of motion.
@@
Numerical methods to identify fold bifurcations mostly consist of pseudo-arclength algorithms coupled with shooting techniques or collocation-based methods. By varying the bifurcation parameter, one of the two branches of periodic solutions is continued until the fold is reached [36]. Alternatively, through direct numerical integration of the equations of motion and by performing a sweep-up and a sweep-down of the bifurcation parameter, the fold is defined by a sudden jump-down and settling of the periodic solution onto an equilibrium (for a bifurcation scenario similar to the one depicted in Fig. 1). Although these methods typically exploit the system’s equations of motion, they can also be implemented through a black-box model of the system.
@@
Similarly, experimental methods for fold bifurcation identification usually implement a sweep-up and a sweep-down of the bifurcation parameter while the system operates. The fold is identified by a sudden jump-down of a periodic solution. Additionally, control-based continuation techniques enable tracking unstable solutions until a fold is reached or even track folds in the parameter space [37].
@@
Several methods to predict the presence of fold bifurcations directly from time series, without actually reaching it, were proposed in the literature. Suppose the system is assumed on the branch of periodic solutions leading to the fold. In that case, these methods usually exploit early-warning signals related to statistical properties of the system’s response to small perturbations [38, 21]. These methods can recognize if a system is close to a fold (in this context, usually referred to as tipping point [39] or critical transition [40, 41]), possibly enabling operators to control the system and avoid reaching the fold, which would be a non-reversible transition.
@@
If the system in operation is on the branch of solutions not directly related to the fold, small perturbations can hardly be used to predict the fold; in fact, the fold has no local effect on the solution. However, large perturbations can still be used to forecast the fold without actually reaching it. Epureanu and coworkers developed similar methods in a series of papers [6, 42, 43, 44]; the methods they developed allow one to predict a full bifurcation diagram from the analysis of the displacement decay of a few trajectories obtained in the pre-bifurcation scenario. A similar technique, specifically targeting fold bifurcations, was recently developed in [45] and tested experimentally in [46].
@@
In this study, we try to perform a similar task utilizing machine learning. In particular, we utilize so-called scientific convolutional neural networks (NNs). NNs have been long used to analyze time series and identify anomalies. Examples include but are not limited to, human activity recognition [47, 48], electrocardiogram classification [49, 50], speech recognition [51, 52], natural language processing [53, 54]. NNs have the ability to handle long-term dependencies and temporal context [55]. They can recognize patterns and are good at performing classifications [56]. All these properties make them a potentially useful tool for identifying if a time series is near a fold bifurcation in the parameter space. Many studies demonstrated the ability of NNs to classify and process time series related to dynamical systems. For example, they are used for chatter detection in machining [57, 58], chaos identification [59], fault detection and diagnosis in mechanical systems [60, 61], anomaly detection [62], and many other similar tasks.
Recently, NNs have also been implemented for bifurcation predictions. For example, in [63], a sort of convolutional NN (CNN), combined with a long-short term memory (LSTM) architecture, is exploited to predict tipping points. The network, trained only with synthetic data from several systems, could predict tipping points from real-life data. A similar result is illustrated in [64], applied to climate systems. In [65], the synchronization of chaotic systems is predicted through NNs. Other studies addressed the problems of bifurcation prediction through NNs in more general terms. For instance, in [66], deep learning is exploited to predict a large range of different bifurcations, training a CNN-LSTM with synthetic data from differential equations and using it on real-life ones. In [67], a similar result is obtained using discrete time equations, also providing promising results. Both studies [66, 67] required using a large number of randomly generated models (in the order of tens of thousands).
@@
Although NNs are excellent interpolators, being able to deal with large inputs, they are generally unable to extrapolate results, meaning that if an input is outside of the range of the input used for training, a NN tends to provide a wrong output [68]. One method to overcome, or at least limit, this problem is to design and train the network, taking into account the physics of the problem as well, and not only data; in other words, by providing the network with some physics-based information. Such NNs are usually referred to as scientific or physics-informed NNs. The main idea behind them is to introduce an appropriate inductive, learning, or observational bias that can steer the learning process towards identifying physically consistent solutions [69].
@@
Inductive bias consists of tailoring the NN architecture in such a way that it implicitly satisfies some physical laws; this approach is usually implemented through some sort of invariant symmetries and rotations; this is done, for instance, by CNNs [70], providing excellent results in computer vision and many other fields.
@@
Learning bias is usually provided through the loss function, which also includes a penalty if a given mathematical function is not satisfied. This mathematical function can have any desired form, and it should represent a physical principle that we want the NN to approximately satisfy, such as, for example, conservation of mass or energy in a mechanical system [71].
@@
Observational bias consists of providing the network with data embodying the underlying physics of the problem or using specifically designed data augmentation procedures. For instance, an observational bias can be obtained by conveniently balancing experimental and synthetic data. This method was successfully applied in a variety of fields, such as, for example, fracture propagation [72].
@@
In this work, we investigate the ability of a scientific CNN to identify trajectories that are close to a fold bifurcation. The NN is physics-informed through an observational bias obtained through an appropriate input normalization, which hides some signal properties (the frequency) and highlights other properties (the oscillation amplitude trend). Several studies in the literature have already approached the same problem through machine learning [66, 73]; however, up to the authors’ knowledge, it is the first time that the prediction is performed from trajectories not converging to the branch related to the fold itself. Our results show that such an NN can classify trajectories with excellent accuracy and possesses remarkable extrapolation properties.
@@ 2 Objective
This study aims to train a CNN to identify if a time series belongs to a parameter set near a fold bifurcation. This objective is formalized in a classification task, as detailed below. The developed CNNs will be trained to distinguish between trajectories referring to a point in the parameter space that is far from a fold (region A in Fig. 1), close to a fold (region B), or after a fold (region C). Considering that the fold bifurcation generates two branches of periodic solutions, regions of the parameter space where the branches of periodic solutions do not exist are interpreted as before the fold (either far from it or close to it). In contrast, regions where they exist are considered as after. The bifurcation parameter is constant for each trajectory.
We consider trajectories that are always initiated from relatively large initial conditions, such that they capture the transient dynamics related to the fold (although it might not be the case, as illustrated in [45]). Other methods for forecasting fold bifurcations have, in fact, the same requirement [42, 45].
@@
All utilized trajectories relative to the bistable region converged to the periodic solution because of their large initial conditions. This was purposely done in order to simplify the computation; in fact, trajectories starting from small amplitude and converging to the equilibrium, even if in the bistable region, can easily be confused with trajectories far from the fold. This simplification does not reflect any specific engineering aspect in terms of application, but it is a necessary step to approach the problem at hand gradually. Future studies will also address small amplitude oscillations, aiming at identifying features distinguishing low amplitude trajectories.
@@
We notice that, while the distinction between regions B and C (close and after the fold) is precisely defined by the bifurcation, the boundary between trajectories far and close to the fold is completely arbitrary. To overcome this problem, trajectories used to train the CNNs of the two groups are very well separated, as explained in Sect. 4.3. Then, the trained CNNs are tested on a large range of values of the bifurcation parameter in order to evaluate their capabilities.
@@ 3.1 Oscillator with nonlinear damping
For $c_{1}>0$, the trivial solution is always stable. For $c_{3}=0$, the system is linear, while for $c_{3}\neq 0$ the damping has a nonlinear characteristic. If $c_{3}>0$, for increasing oscillation amplitudes, first damping decreases and then increases because of the cubic and fifth order terms. As thoroughly investigated in [28], for $c_{3}=c_{3}^{*}:=40c_{1}/9$ the system presents a fold bifurcation, which generates two branches of periodic solutions existing for $c_{3}>c_{3}^{*}$. Increasing $c_{3}$, the branch of unstable periodic solutions approaches the trivial solution, reducing its dynamical integrity. However, as far as $c_{1}>0$, they never merge. The corresponding bifurcation diagram is illustrated in Fig. 2(a).
@@
Figure 2: Bifurcation diagrams of the investigated systems; solid lines: stable solutions, dashed lines: unstable solutions; red dots: fold bifurcations. The lines mark the maximal amplitude of the steady-state solutions for variations of the bifurcation parameter. (a) System in Eq. (1) for $c_{1}=0.5$; (b) system in Eq. (2); (c) system in Eq. (5) for the parameter values $r=0.05$, $\gamma=0.97$, $\mu_{2}=0.12$ and $\alpha=0.3$; (d) system in Eq. (7).